the wavelength of a certain radiation is 6625 angstroms then the number of quantas present in 1J of energy of that radiation is

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Dear student 
From Planck’s Equation energy of a photon
E=hc/l
Energy of 1 photon = (6.62 × 10^-34 )×(3 × 10^8)/(6625 × 10^-12)
=>2.9977 * 10^(-17) Joules
SO, 
No, of photons required for 1 J = 1/(2.997 * 10^(-17)
=> 3.3366 * 10^16 Photons